# Boosting quantum annealing performance through direct polynomial unconstrained binary optimization

### v2

@@ Abstract
Quantum annealing aims at solving optimization problems of practical relevance using quantum-computing hardware. Problems of interest are typically formulated in terms of quadratic unconstrained binary optimization (QUBO) Hamiltonians. However, many optimization problems are much more naturally formulated in terms of polynomial unconstrained binary optimization (PUBO) functions of higher order. As we show with various problem examples, leveraging the PUBO formulation can bring considerable savings in terms of required number of qubits. Moreover, in numerical benchmarks for the paradigmatic 3-SAT problem, we find scenarios where the scaling of the minimal gap during the optimization sweep differs significantly, suggesting the possibility of an exponentially faster annealing time when using the PUBO as compared to the QUBO formulation. This advantage persists even when considering the overhead in implementing the higher-order interactions necessary for PUBO cost Hamiltonians. As an interesting side effect, the analysis on minimum energy gaps of different 3-SAT instance generators reveals different degrees of hardness, which will be of interest also for classical benchmark calculations. Our findings show a promising path to improving the resource efficiency and sweeping speed of quantum annealing, important prerequisites when aiming at solving larger optimization problems with relevance to industry.
@@ I Introduction
The continuous improvements of quantum computing devices have positioned quantum annealing as a promising candidate for tackling combinatorial optimization problems of practical relevance [1, 2, 3, 4, 5]. Experiments on commercially available quantum annealers have recently even claimed quantum advantage [6]. Since most of the existing quantum computers consist of discrete two-level systems coupled by two-qubit gates, the most common approach for solving combinatorial optimization problems on quantum devices is via quadratic unconstrained binary optimization (QUBO), where the cost function is a polynomial of degree two. However, many problems are more naturally expressed as polynomials of higher degrees, i.e., in terms of polynomial unconstrained binary optimization (PUBO). Application-relevant examples of PUBO problems range from vehicle routing [7] to polymer sampling [8, 9] and lattice protein folding [10, 11]. To implement these on quantum hardware, the PUBO problems are usually reduced to QUBO problems [3, 12], and efforts have been made to render the reduction of PUBO to QUBO instances as efficient as possible [13, 14]. Nevertheless, the QUBO reduction comes at the price of introducing additional ancillary variables, along with constraints that enforce their consistency, both of which is best avoided considering the stringent space and interaction-scale limitations of current NISQ devices. Then again, considering the overhead in engineering the associated multi-qubit interactions, it remains an open question whether a direct PUBO implementation could be advantageous in quantum optimization [15].
@@
Here, we answer this question in the affirmative, by highlighting the potential for a gain in efficiency through a direct implementation of PUBO problems. Paradigmatic problems that are naturally formulated in terms of PUBO range from the Boolean satisfiability problem (SAT), over the hypergraph coloring problem, to the $p$-spin model. Computing the number of qubits necessary to implement such problems as QUBO instances, we find a potential resource saving of up to an order of magnitude in number of qubits for some problems when formulating them instead in terms of PUBO problems. Additionally, we compute the minimum energy gap across a range of problems by employing exact diagonalization on small-scale systems, specifically 3-SAT instances of varying complexity generated using different methods [16]. We find the direct formulations as PUBO instances to have larger gaps than the equivalent QUBO reductions, meaning the employment of PUBO formulations can signify a considerable improvement of computing time in practical applications. Finally, we discuss the possibility to synthesize the necessary 3-body spin interactions in quantum computers with typical gate sets employing single-qubit gates combined with CNOT or $R_{zz}$ gates. The required overhead of only four two-qubit gates is more than compensated by the increased gap, in addition to the saving of ancillary qubits. As our results illustrate, it can be of large practical advantage to directly implement PUBO formulations in quantum annealing protocols, thus increasing the scalability of these machines.
@@
On top of that, our studies on the scaling of the minimum energy gap also shed light on the hardness of instances created by various 3-SAT benchmark generators [16]. Among the tested generators, we find completely randomly generated instances to be on average easier to solve (larger minimum energy gap) than the structured instances with fewer logical clauses constructed by one of the benchmark generators. In contrast, for yet another benchmark generator, the formulation as a PUBO shows that it generates only trivially solvable instances, reflected in a large energy gap. Thus, our analysis on the paradigmatic 3-SAT optimization problem, carried out with quantum computers in mind, can also inform benchmark calculations on classical computers.
@@
This paper is organized as follows. Section II contains a brief recap on solving combinatorial optimization problems with quantum annealing and its main limitation due to the closing of the minimum energy gap. The behavior of the minimum gap will be our main figure of merit for assessing potential performance gains of PUBO over QUBO. In section III, we present a series of optimization problems that are naturally formulated as PUBO. We also discuss their reduced spatial resource requirements (i.e., number of qubits) as compared to equivalent QUBO reductions. Section IV puts forward our numerical results on the behavior of the minimum energy gap at the example of the paradigmatic 3-SAT problem, illustrating the potential for PUBO speedups due to larger gaps. In section V, we discuss the synthesis of three-body interactions using single- and two-qubit gates, and compare the resulting overhead with the expected performance gain. Finally, section VI contains our conclusions on the potential advantages of PUBO over QUBO in terms of spatial and temporal resources.
@@ II.3 Estimating PUBO speedup for quantum annealing
III Selected PUBO problems and their resource requirements
In this section, we list a number of optimization problems that are naturally formulated in terms of PUBO instances. We put a focus on cubic problems because the corresponding three-body Hamiltonians typically represent the least complex beyond-quadratic interactions that can be implemented on present-day quantum devices.
We start in section III.1 by illustrating the advantages of the PUBO formulation in terms of spatial resources, i.e., number of required qubits, for a simple toy model of minimizing a polynomial function. In section III.2, we dive in some detail into the paradigmatic 3-SAT problem, which is the model we use for our numerical analysis further below. Section III.3 briefly presents a collection of further examples, illustrating the ubiquity of PUBO problems.
@@ III.1 Toy model: minimization of a polynomial function
As a first simple example, we consider a continuous polynomial $f:]a,b[\to\mathbb{R}$ as an objective function to be minimized. Without loss of generality, we choose the boundaries of the open interval as $a=0$ and $b=2^{n}$ for a convenient encoding of the continuous variable into bits. We can then represent any continuous variable $x$ approximately as a bitstring $(x_{1},\dots,x_{n+m})$ via the mapping
@@
As a concrete case, take the task to find the minimum of the cubic polynomial $f(x)=x^{3}+x$ in the domain $]0,2^{2}[$. After inserting the encoding in eq. 9 with a resolution of $m=1$, we obtain
@@
Due to the term $\propto x_{1}x_{2}x_{3}$, this optimization problem is of PUBO form.
Equation 10 can be reduced to a QUBO problem by introducing an additional slack variable $y=x_{2}x_{3}$. An appropriate penalty term $\propto(3y+x_{2}x_{3}-2x_{2}y-2x_{3}y)$ needs to be added to the optimization problem in order to enforce consistency. While in this specific case the QUBO formulation requires only a single extra qubit compared to the equivalent PUBO formulation, the resource savings can get significant for larger and more complex problems (see below). Furthermore, the optimal introduction of slack variables (i.e., the assignment which requires the least amount of additional bits) is itself known to be an $\mathsf{NP}$-hard problem [25]. Finally, the constraint that enforces consistency of the ancillary variables needs to be sufficiently strong such that other terms cannot overpower it. Since the maximal energy scales available in a given quantum machine are limited, this means in practice reducing the strength of the other terms, and thus incurring proportionally longer sweep times or deteriorating the adiabaticity of the sweep.
@@ III.2 Boolean satisfiability problem
3-SAT can be reformulated in various ways as a PUBO or QUBO problem in order to make it suitable for quantum annealing. A straightforward encoding consists in a PUBO of degree three: each clause is cast into the form of an energy term penalizing any choice of truth values that does not satisfy the clause. For example, the clause $C_{m}=x_{i}\lor\lnot{x_{j}}\lor\lnot{x_{\ell}}$ from above can be translated to $(1-x_{i})x_{j}x_{\ell}$, which only evaluates to zero if at least one variable satisfies the corresponding clause. Adding up these terms for each clause and translating the binary variables to quantum spin variables via $x_{i}=1/2-\hat{s}^{z}_{i}$, leads to the general three-body spin Hamiltonian in eq. 3. Note that, in this specific PUBO formulation, the three-body and two-body interaction strengths $J^{(3)}$ and $J^{(2)}$ as well as the local biases $\bm{h}^{z}$ are all of the same order of magnitude or zero. The ground state of the cost Hamiltonian then encodes the solution of the original 3-SAT problem (more precisely, it solves MAX-SAT, i.e., finding the solution that satisfies most clauses). This encoding requires $N$ spins (qubits) for the $N$ Boolean variables in the original problem.
@@
III.2.3 3-SAT as QUBO via the maximum independent set problem
Finding the optimal QUBO reduction of 3-SAT is by itself an $\mathsf{NP}$-hard problem. However, other non-optimal generic QUBO formulations of 3-SAT are possible. A well-known example is the formulation of 3-SAT as the equivalent maximal independent set problem (MIS) [37]: Draw a graph with three vertices for each clause (corresponding to the three involved literals). All vertices belonging to the same clause as well as pairs of vertices corresponding to mutually exclusive literals (i.e., $x_{i}$ and $\lnot{x_{i}}$) are connected by edges, the set of which is denoted with $E$. The MIS problem then asks whether it is possible to color $M$ vertices without two colored vertices being connected by an edge. If the answer is yes, the original 3-SAT problem is satisfiable. This MIS problem can be easily translated to a QUBO problem requiring $3M$ qubits (or about $\operatorname*{round}(3\times 4.24\,N)=\operatorname*{round}(12.72\,N)$ for hard problems):
$f_{\mathrm{MIS}}(\bm{x})=\sum\limits_{ij\in E}x_{i}x_{j}\,.$ (13)
Although generically more resources are required than for the reduction with slack variables, one avoids the $\mathsf{NP}$-hard problem of finding an efficient assignment of slack variables. Another potential advantage of the MIS formulation, especially regarding its use on physical quantum devices, lies in the fact that all interactions have the same energy scale, whereas slack variables for a direct QUBO reduction are introduced with constraints that are imposed through high energy penalties.
III.2.4 3-SAT as QUBO via linear inequality constraints
$\text{max}\sum_{m}z_{m}$ (14)
under the inequality constraints
III.3 Further PUBO problems
There exists a plethora of further optimization problems that are naturally posed as PUBO. Below, we mention a few relevant examples.
III.3.1 $p$-spin model
The cost Hamiltonian of the $p$-spin model ($p\geq 2$) reads
III.3.2 Hypergraph coloring
III.3.3 NAE-k-SAT
Apart from $k$-SAT, which we have discussed in detail above, there are many other variations and constraints on the general satisfiability problem. One of those is the NAE-$k$-SAT (Not-All-Equal-k-SAT) problem [49], which can also be naturally formulated as a PUBO. Like $k$-SAT, NAE-$k$-SAT consists of a conjunction of clauses, where each clause contains $k$ Boolean variables and is only true, if at least one variable is true and at least one variable is false. Only for $k\geq 4$ is the problem naturally formulated as a PUBO with higher-order interactions [49]. For example, for $k=4$, a clause $C(x_{i},x_{j},x_{k},x_{l})$ can be encoded into one four-body interaction $\sim\hat{s}^{z}_{i}\hat{s}^{z}_{j}\hat{s}^{z}_{k}\hat{s}^{z}_{l}$ and six two-body interaction penalty terms of the cost Hamiltonian.
III.3.4 Traveling salesperson problem with time windows
@@ IV.1.1 Characteristic energy scales
In our numerics, we express the cost Hamiltonian $\hat{H}_{\mathrm{cost}}$ in units of the characteristic energy scale $J$. In practice, $J$ is limited by the maximum physical interaction rate (or inverse gate time) available on the quantum device. Usually, interactions of higher rank are slower compared to interactions of lower rank. Therefore, the characteristic energy scale for cost Hamiltonians encoding PUBO problems of third order is determined by the physical three-body interaction strength (inverse three-qubit gate time), while for QUBO problems, the two-body interaction rate (inverse two-qubit gate time) is relevant. To account for this, we renormalize the cost Hamiltonian in eq. 3 such that $J=\max_{ijk}|J_{ijk}^{(3)}|$ for PUBO and $J=\max_{ij}|J_{ij}^{(2)}|$ for QUBO.

### v3

# thanks: Corresponding authorthanks: Corresponding author

@@ Abstract
Quantum annealing aims at solving optimization problems of practical relevance using quantum-computing hardware. Problems of interest are typically formulated in terms of quadratic unconstrained binary optimization (QUBO) Hamiltonians. However, many optimization problems are much more naturally formulated in terms of polynomial unconstrained binary optimization (PUBO) functions of higher order. As we show with various problem examples, leveraging the PUBO formulation can bring considerable savings in terms of required number of qubits. Moreover, in numerical benchmarks for the paradigmatic 3-SAT problem, we find scenarios where the scaling of the minimum energy gap during the optimization sweep differs significantly, suggesting the possibility of an exponentially faster annealing time when using the PUBO as compared to the QUBO formulation. This advantage persists even when considering the overhead caused by the higher-order interactions necessary for PUBO cost Hamiltonians. As an interesting side effect, the analysis on minimum energy gaps of different 3-SAT instance generators reveals different degrees of hardness, which will be of interest also for classical benchmark calculations. Our findings show a promising path to improving the resource efficiency and sweeping speed of quantum annealing protocols on both analog and digital platforms, which are important prerequisites when aiming at solving larger optimization problems with relevance to industry.
@@ I Introduction
The continuous improvements of quantum computing devices have positioned quantum annealing as a promising candidate for tackling combinatorial optimization problems of practical relevance [1, 2, 3, 4, 5]. Experiments on commercially available quantum annealers have recently even claimed quantum advantage [6]. Since most of the existing quantum computers consist of discrete two-level systems coupled by two-qubit gates, the most common approach for solving combinatorial optimization problems on quantum devices is via quadratic unconstrained binary optimization (QUBO), where the cost function is a polynomial of degree two. However, many problems are more naturally expressed as polynomials of higher degrees, i.e., in terms of polynomial unconstrained binary optimization (PUBO), sometimes also referred to as higher-order binary optimization (HOBO). Application-relevant examples of PUBO problems range from vehicle routing [7] to polymer sampling [8, 9] and lattice protein folding [10, 11]. To implement these on quantum hardware, the PUBO problems are usually reduced to QUBO problems [3, 12], and efforts have been made to render the reduction of PUBO to QUBO instances as efficient as possible [13, 14]. Nevertheless, the QUBO reduction comes at the price of introducing additional ancillary variables, along with constraints that enforce their consistency, both of which is best avoided considering the stringent space and interaction-scale limitations of current NISQ devices. Then again, considering the overhead incurred by the associated multi-qubit interactions, i.e., slower interaction rates on analog platforms or increased circuit depth on digital architectures, it remains an open question whether a direct PUBO implementation could be advantageous in quantum optimization [15].
@@
Here, we answer this question in the affirmative, by highlighting the potential for a gain in efficiency through a direct implementation of PUBO problems. Paradigmatic problems that are naturally formulated in terms of PUBO range from the Boolean satisfiability problem (SAT), over the hypergraph coloring problem, to the $p$-spin model. Computing the number of qubits necessary to implement such problems as QUBO instances, we find a potential resource saving of up to an order of magnitude in the number of qubits for some problems when formulating them in terms of PUBO problems instead. Additionally, we compute the minimum energy gap across a range of problems by employing exact diagonalization on small-scale systems, specifically 3-SAT instances of varying complexity generated using different methods [16]. We find the direct formulations as PUBO instances to have larger gaps than the equivalent QUBO reductions, suggesting an improvement in computing time. While for hard instances the gap closing and thus the time-to-solution nonetheless scale exponentially with the problem size, direct PUBO formulations can exhibit smaller exponents for some classes of 3-SAT problems, which translates to an exponential speedup over the corresponding QUBO formulation. This behavior is agnostic to the underlying quantum computing architecture and holds for analog as well as digital platforms alike.
For gate-based quantum computers, we discuss the possibility to synthesize the necessary 3-body spin interactions with typical gate sets employing single-qubit gates combined with CNOT or $R_{zz}$ gates. The required overhead of only four two-qubit gates per three-qubit gate is more than compensated by the saving of ancillary qubits as well as the increased minimum energy gap. This shows that it can be of large practical advantage to directly implement PUBO formulations in quantum annealing protocols, thus increasing the scalability of these machines.
@@
On top of that, our studies on the scaling of the minimum energy gap also shed light on the hardness of instances created by various 3-SAT benchmark generators [16]. Among the tested generators, we find completely randomly generated instances to be on average easier to solve (larger minimum energy gap) than the structured instances with fewer logical clauses constructed by one of the benchmark generators. In contrast, for yet another benchmark generator, the formulation as a PUBO shows that it generates only trivially solvable instances, reflected in a large energy gap. Thus, our analysis of solving 3-SAT on quantum computers can also inform benchmark calculations for combinatorial optimization problems on classical computers, where recent studies have demonstrated significant performance gains through quantum-inspired metaheuristic solvers for problems as large as 100 000100000100\,000100 000 variables [17, 18, 19].
This paper is organized as follows. Section II contains a brief recap on solving combinatorial optimization problems with quantum annealing and its main limitation due to the closing of the minimum energy gap. The behavior of the minimum gap will be our main figure of merit for assessing potential performance gains of PUBO over QUBO.
@@
In section III, we discuss the reduced spatial resource requirements (i.e., number of qubits) of optimization problems that are naturally formulated as PUBO, as compared to equivalent QUBO reductions. Section IV puts forward our numerical results on the behavior of the minimum energy gap with the example of the paradigmatic 3-SAT problem, illustrating the potential for PUBO speedups due to larger gaps. In section V, we discuss the synthesis of three-body interactions using single- and two-qubit gates, and compare the resulting overhead with the expected performance gain. Finally, section VI contains our conclusions on the potential advantages of PUBO over QUBO in terms of spatial and temporal resources.
@@ II.3 Estimating PUBO speedup for quantum annealing
II.4 Analog versus digital quantum annealing
Quantum annealing has originally been conceptualized for analog platforms, where the relative strength of the cost and driving Hamiltonian can be tuned continuously. With the increasing availability of digital quantum computers, variants of quantum annealing adapted to gate-based architectures have become popular. Digital annealers are particularly interesting in the light of direct implementations of PUBO as they allow one to synthesize the required higher-order interactions from universal sets of two- and single-qubit gates, whereas engineering native higher-order interactions in analog annealers is more challenging (see discussion in section V.1). For this reason, we briefly introduce a digital version of quantum annealing in this section. Our analysis of PUBO versus QUBO performance applies to both analog and digital protocols alike as long as they are based on adiabatic evolution (in distinction to certain quantum-annealing-inspired algorithms like the Quantum Approximate Optimization Algorithm (QAOA) [28], where optimal performance requires schedules with large angles corresponding to evolution times outside the adiabatic regime).
Digital quantum annealing (or Trotterized quantum annealing) is accomplished via a Trotterization of the time evolution under the Hamiltonian in eq. 4. That is, for a grid of times $0=t_{0}<t_{1}<\dots<t_{M}=t_{\mathrm{sweep}}$ with uniform step size $\Delta t=t_{\mathrm{sweep}}/M$, the time evolution operator can be approximated as [29]
$\hat{\mathcal{T}}\mathrm{e}^{-it_{\mathrm{sweep}}\int_{0}^{1}\mathop{}\!\mathrm{d}s^{\prime}\,\hat{H}(s^{\prime})/\hbar}\\ =\prod_{m=1}^{M}\mathrm{e}^{-i\Delta t(1-s_{m})\hat{H}_{\mathrm{drive}}/\hbar}\mathrm{e}^{-i\Delta ts_{m}\hat{H}_{\mathrm{cost}}/\hbar}+\mathcal{O}(\Delta t)\,,$ (9)
where $\hat{\mathcal{T}}$ is the time-ordering operator and $s_{m}=t_{m}/t_{\mathrm{sweep}}$. Note that the local error for each Trotter step is $\mathcal{O}(\Delta t^{2})$, such that the accumulated global error scales as $\mathcal{O}(\Delta t)$, which may be improved through higher-order splitting methods [30]. Since all terms within the cost Hamiltonian commute with each other, the corresponding matrix exponential can be exactly decomposed into a series of single-qubit $Z$ rotations and two-qubit $ZZ$ rotations. Similarly, the matrix exponential of the driving Hamiltonian can be decomposed into distinct single-qubit $X$ rotations. Each two-qubit rotation may then further be compiled into the native gate set of the respective architecture.
Compared to analog quantum annealing, this procedure incurs additional errors due to the discretization. Furthermore, since each time step $\Delta t$ needs to be decomposed into the native gate set, there is in general a polynomial overhead in the number of gates. For sufficiently small Trotter steps, however, the limiting factors to the algorithm’s performance are the same as those of the continuous protocol. In this situation, the adiabaticity time estimated from the minimum energy gap of the analog quantum annealing protocol acts as a proxy for the number of discrete time steps and thus the circuit depth required to reach the ground state of the cost Hamiltonian with the discrete protocol. Unless stated otherwise, we will generally consider analog quantum annealing throughout this work.
III PUBO problems and their resource requirements
In this section, we discuss two instructive examples of optimization problems that can naturally be formulated as PUBO. We start in section III.1 by illustrating the advantages of the PUBO formulation in terms of spatial resources, i.e., number of required qubits, for the simple toy model of minimizing a polynomial function. In section III.2, we dive in some detail into the paradigmatic 3-SAT problem, which is the model we use for our numerical analysis in the rest of this work.
Appendix A briefly presents a collection of further examples (the $p$-spin model, hypergraph coloring, NAE-k-SAT, and the traveling salesperson problem with time windows), illustrating the ubiquity of PUBO problems.
@@ III.1 Toy model: minimization of a polynomial function
As a first simple example, we consider a continuous polynomial $f:[a,b)\to\mathbb{R}$ as an objective function to be minimized. Without loss of generality, we choose the boundaries of the half-open interval as $a=0$ and $b=2^{n}$ for a convenient encoding of the continuous variable into bits. We can then represent any continuous variable $x$ approximately as a bitstring $(x_{1},\dots,x_{n+m})$ via the mapping
@@
As a concrete case, take the task to find the minimum of the cubic polynomial $f(x)=x^{3}+x$ in the domain $[0,2^{2})$. After inserting the encoding in eq. 10 with a resolution of $m=1$, we obtain
@@
Due to the term $\propto x_{1}x_{2}x_{3}$, this optimization problem is of third-order PUBO form. The trivial solution in this case is given by $(x_{1},x_{2},x_{3})=(0,0,0)$.
Equation 11 can be reduced to a QUBO problem by introducing an additional slack variable $y=x_{2}x_{3}$. An appropriate penalty term $\propto(3y+x_{2}x_{3}-2x_{2}y-2x_{3}y)$ needs to be added to the optimization problem in order to enforce consistency. While in this specific case the QUBO formulation requires only a single extra qubit compared to the equivalent PUBO formulation, the resource savings can get significant for larger and more complex problems (see below). Furthermore, the optimal introduction of slack variables (i.e., the assignment which requires the least amount of additional bits) is itself known to be an $\mathsf{NP}$-hard problem [31]. Finally, the constraint that enforces consistency of the ancillary variables needs to be sufficiently strong such that other terms cannot overpower it. Since the maximal energy scales available in a given quantum machine are limited, this means in practice reducing the relative strength of the other terms, and thus incurring proportionally longer sweep times or deteriorating the adiabaticity of the sweep. Finding the minimum penalty strength that enforces all constraints is in general $\mathsf{NP}$-hard, but efficient approximate schemes with significant performance gains over default choices have been proposed [32].
@@ III.2 Boolean satisfiability problem
3-SAT can be reformulated in various ways as a PUBO or QUBO problem in order to make it suitable for quantum annealing. A straightforward encoding consists in a PUBO problem of degree three: each clause is cast into the form of an energy term penalizing any choice of truth values that does not satisfy the clause. For example, the clause $C_{m}=x_{i}\lor\lnot{x_{j}}\lor\lnot{x_{\ell}}$ from above can be translated to $(1-x_{i})x_{j}x_{\ell}$, which only evaluates to zero if at least one variable satisfies the corresponding clause. Adding up these terms for each clause and translating the binary variables to quantum spin variables via $x_{i}=1/2-\hat{s}^{z}_{i}$, leads to the general three-body spin Hamiltonian in eq. 3. Note that, in this specific PUBO formulation, the three-body and two-body interaction strengths $J^{(3)}$ and $J^{(2)}$ as well as the local biases $\bm{h}^{z}$ are all of the same order of magnitude or zero. The ground state of the cost Hamiltonian then encodes the solution of the original 3-SAT problem (more precisely, it solves MAX-SAT, i.e., finding the solution that satisfies most clauses). This encoding requires $N$ spins (qubits) for the $N$ Boolean variables in the original problem.
@@
In appendix B, we discuss two other approaches to formulate 3-SAT as a QUBO problem. However, the direct approach of introducing slack variables is the most efficient in terms of qubit requirements and will be considered for the rest of this work.
@@ IV.1.1 Characteristic energy scales
In our numerics, we express the cost Hamiltonian $\hat{H}_{\mathrm{cost}}$ in units of the global energy scale $J$, which we identify with a characteristic energy of the physical device. In principle, it is desirable for $J$ to be as large as possible in order to speed up the absolute computing time. In practice, however, $J$ is limited by the maximum physical coupling rates the device can operate with. Consequently, a global rescaling of the Hamiltonian is required to ensure that none of the terms in eq. 3 exceed these limits. On analog hardware, the physical couplings relevant for PUBO implementations up to third order include the maximum physical three-body and two-body interaction strengths as well as the maximum strength of local fields. On digital devices, the physical time scales associated with these interactions correspond, respectively, to the three-, two-, and single-qubit gate times. Since interactions of higher rank are generally slower compared to interactions of lower rank (see discussion in section V.1), the former typically create a bottleneck: the global energy scale $J$ is limited by the maximum physical three-body interaction strength if three-body interactions are present (third-order PUBO) and by the maximum physical two-body interaction strength if only two-body interactions are present (QUBO), where we assume that local fields can be scaled sufficiently high to pose no restriction. In particular, three-body interactions are the limiting factor for PUBO implementations of typical 3-SAT instances encountered in our small-scale numerical experiments, where the maximum ratio of three-body and two-body couplings is of order one (though this may be different for larger instances or problems with different PUBO structures). It is therefore natural in our case to choose the characteristic energy $J=\max_{ijk}|J_{ijk}^{(3)}|$ as the global energy unit for third-order PUBO, while for QUBO we use $J=\max_{ij}|J_{ij}^{(2)}|$.